Q.1 Consider the following information about a hypothetical economy:

1. Y = A(0.025K-0.5N)N

2. A=2/3

3. K=2000

4. N^s=-18+(18/5)w

5. C=200+(2/3)(Y-T)-300r

6. T=-75+(1/4)Y

7. I =100−100r

8. G =100

9. L = 0.5Y − 200i

10. M = 6300

11. π^e= 0.10

Now using this information, answer the following:

(i) Starting from the initial equilibrium position again, suppose that the capital stock increases by 170. What 

will be the impact of this expansion on labour market equilibrium and aggregate supply of output? 

Calculate values of all endogenous variables and give intuitive explanation of the results. 

(j) Compare the equilibrium positions in (d) and (i) indicating all points. 

(k) Suppose that Li → ∞ in Equation.9 of the model. How will it affect the shape of the money demand and 

the LM curve. Will the monetary policy of part (e) have the same effect as calculated above or any 

different? Explain using graphs and multipliers.


1
Expert's answer
2021-07-05T08:58:31-0400

a

"Y=\frac{2}{3}[(0.025×2000)-0.5N]N\\Y=\frac{2}{3}(50-0.5N)N\\Y=\frac{100N}{3}-\frac{1}{3}N^2"

equilibrium income


"T=-75+\frac{1}{4}Y\\T=-75+\frac{25y}{3}-\frac{y^3}{12}"

taxes


"C=300+\frac{2}{3}(Y-T)-300r\\C=300+\frac{2}{3}(Y-(-75+\frac{1}{4}y))-300r\\C=-300r+\frac{y}{2}+250\\C=-300r+250+\frac{\frac{100N}{3}-\frac{N^3}{3}}{2}"

consumption


"N=-18+\frac{18}{5}W\\Y=\frac{594}{5}W-594\\Y-594=\frac{594}{5}W\\W=\frac{5(Y-594)}{594}\\N=-18+\frac{18}{5}(\frac{5(Y-594)}{594})\\N=\frac{Y-1188}{33}"


"I=100-100r\\r=\frac{1000-r}{100}\\interest \space rates\\"


any change in in values affects each other because they are corelated. the economy depends on several factors to be at equilibrium.


b

the endogenous variables include Y

and exogenous variables C, I,T,C